OpenStudy (elcna):

Can someone check my answers please? :) 1. Solve the equation using square roots. x^2-81=0 *-√9,√9 <<< *-81,81 *-9,9 *no real number solutions 2. 3x^2-147=0 *-49,49 *-7,7 *-√7,√7 <<< *no real number solutions 3. The area of a playground is 20 square yards. The length of the playground is 5 times longer than its width. Find the length and width of the playground. *length=1 yd, width=20 yd *length=2 yd, width=10 yd *length=10 yd, width=2 yd *length=20 yd, width=1 yd For this one I was kind of caught between B and D. Any help is appreciated!

hartnn (hartnn):

\(x^2 = 81 \\ x^2= 9^2 \\ x = \pm 9\) and not square root of 9. same error for 2nd

OpenStudy (elcna):

Thank you so much! @hartnn

hartnn (hartnn):

still need help for 3rd?

OpenStudy (elcna):

If you could, that'd be great :)

hartnn (hartnn):

let the width be x length = 5x Area = length*width = 5x^2 = 20 get x from here :)

OpenStudy (elcna):

Thank you!!

hartnn (hartnn):

np, bdw what did you get for 3rd?

hartnn (hartnn):

just to cross-check...

OpenStudy (elcna):

I got 2,-2 ?

hartnn (hartnn):

we discard -2 as width can't be negative

hartnn (hartnn):

so width = 2 length = 10 :)

OpenStudy (elcna):

You're a life saver, thank you so much :)

hartnn (hartnn):

welcome ^_^